# Recent Advances of Foundation Language Models-based Continual Learning: A Survey

### v1

@@ 6.2. Online Datasets for NLP
The foundational text classification benchmark, as introduced by Zhang et al. (Zhang et al., 2015) has traditionally been applied in offline continual learning settings. Recent advancements have adapted this benchmark for online continual learning, notably in studies such as MBPA++ (de Masson D’Autume et al., 2019) and OML-ER (Holla et al., 2020)
@@ 6.3. Offline CL Datasets for Multi-modal Tasks
6.4. Offline and Online Datasets for Other Tasks
In this paper, we provide a comprehensive review of recent studies in natural language processing (NLP) and multi-modal tasks, with a particular focus on the use of PLMs, LLMs and VLMs. Additionally, we introduce a range of offline continual learning (CL) datasets employed across various applications, including automatic speech recognition (Vander Eeckt et al., 2023; Della Libera et al., 2023), autonomous driving (Verwimp et al., 2023), disease classification (Derakhshani et al., 2022), reinforcement learning (Wołczyk et al., 2021), graph data (Ko et al., 2022), and computer vision (Faber et al., 2023; Lin et al., 2021). Furthermore, some online CL benchmarks are also proposed for computer version tasks, such as continual visual learning (Cai et al., 2021), continual object detection (Wang et al., 2021b).
@@ 7. Metrics
Memory stability is typically evaluated by backward transfer (BWT) (Lopez-Paz and Ranzato, 2017) and forgetting measure (FM) (Chaudhry et al., 2018).
Backward Transfer (BWT) emerges as a pivotal concept extensively documented in the literature, notably by Lopez et al. (Lopez-Paz and Ranzato, 2017) and Wu et al. (Wu et al., 2023). BWT measures the performance degradation on previously mastered tasks after the model is trained on new tasks. This performance degradation phenomenon is often referred to as “forgetting”.
@@
Additionally, Chaudhry et al. (Chaudhry et al., 2018) introduce the Forgetting Measure (FM), a metric designed to quantify the extent of forgetting a model experiences for a specific task. A lower FM indicates better retention of previous tasks. Davari et al. (Davari et al., 2022) propose a method named linear probes (LP) to assess representation forgetting. This approach measures the effectiveness of learned representations via an optimal linear classifier trained on the frozen activations of a base network. Representation forgetting is quantified by evaluating the change in Language Processing (LP) performance before and after the introduction of a new task. Kemker et al. (Kemker et al., 2018) introduce three metrics, where $\Omega_{\text{base}}$ assesses retention of initial learning, $\Omega_{\text{new}}$ measures recall of new tasks, and $\Omega_{\text{all}}$ evaluates overall proficiency in maintaining old knowledge and acquiring new information. Additionally, researchers (Koh et al., 2022) devise a novel metric, termed the Knowledge Loss Ratio (KLR), quantifies knowledge degradation using principles from information theory.
@@
Evaluating learning plasticity can be effectively accomplished through two key metrics: forward transfer (FWT) (Lopez-Paz and Ranzato, 2017) and intransigence measure (IM) (Chaudhry et al., 2018).
Forward Transfer (FWT) (Lopez-Paz and Ranzato, 2017) assesses the beneficial effects on the performance of subsequent tasks following a model’s training on prior tasks.
@@
where $R_{0,i}$ denotes the performance metric associated with training on task $i$ independently. Higher values of FWT indicate superior model performance. It is important to note that discussing backward transfer for the initial task is not applicable, as there are no preceding tasks to influence its performance.
Intransigence measure (IM), as defined by Chaudhry et al. (Chaudhry et al., 2018), quantifies a model’s inability to learn new tasks. This measure is calculated by comparing the performance difference of a task when trained jointly with other tasks versus when trained in a continual learning setting. Moreover, Koh et al. (Koh et al., 2022) introduce novel metrics, known as Knowledge Gain Ratio (KGR), which quantifies the capacity to acquire new knowledge through the calculation of knowledge gain.
@@
CKL (Jang et al., 2022b) introduces a novel metric, named FUAR (FORGOTTEN / (UPDATED + ACQUIRED) RATIO), which quantitatively measures the efficiency of each CKL method. It calculates the number of instances of time-invariant knowledge that a model forgets in order to learn or update one instance of new knowledge. When FUAR is equal to 1.0, it signifies an equilibrium where one time-invariant knowledge instance is forgotten on average to obtain a new or updated knowledge instance.
@@
Near-future accuracy (NFA) (Al Kader Hammoud et al., 2023) is introduced as a novel evaluation metric for OCL problem. Unlike traditional evaluation methods that assess models on immediately subsequent samples, NFA evaluates models on samples slightly further into the future, using a minimal shift $S$. Such operation can mitigate label correlation effects, which can adversely impact the accuracy of model adaptability assessments. The smallest shift $S$ is selected to ensure that the test sample aligns closely with the distribution of recently observed training data. Yogatama et al. (Yogatama et al., 2019) proposed a novel online codelength, inspired by prequential encoding (Blier and Ollivier, 2018), to quantify how quickly an existing model can adapt to a new task.
@@ 8. Challenges and Further Work
Most existing studies in the domain of continual learning assume static datasets with known distributions in a relatively closed environment. Moreover, these studies mainly focus on simple tasks (e.g., text classification, sentiment analysis and intent classification) with clear labels. These assumptions do not hold in real-world applications, where environments continually evolve and introduce novel stimuli. A key challenge is to develop continual learning models that operate effectively in complex, noisy environments where clear labels are not always available, and task domains frequently change. Liu et al. (Liu et al., 2023b) recently proposed the SOLA framework to address these limitations by facilitating autonomous adaptation in AI systems. Despite this progress, significant challenges remain in enabling these systems to independently adjust to new, dynamic environments without ongoing human oversight. Future research should focus on developing algorithms capable of autonomously detecting and adapting to shifts in data distribution, thereby improving the applicability of AI in dynamic real-world scenarios.
@@
Privacy protection in continual learning systems poses a significant challenge, particularly as these systems are designed to continuously update and refine their models based on incoming data streams. Unlike traditional static machine learning models, continual learning systems frequently access and process sensitive data across different contexts and time periods, raising substantial concerns about data confidentiality and user privacy. Effective privacy-preserving mechanisms must be integrated into the architecture of these systems to ensure that they do not inadvertently expose or misuse personal data. Techniques such as differential privacy (Dwork, 2006), federated learning (Zhang et al., 2021), and secure multi-party computation (Goldreich, 1998) offer promising solutions by allowing models to learn from decentralized data sources without needing to access the actual data directly. Future research in continual learning should not only focus on enhancing learning efficiency and adaptability but also prioritize the development of robust frameworks that safeguard user privacy across all phases of data handling and model updating.

### v2

@@ 7.2. Online Datasets for NLP
The foundational text classification benchmark, as introduced by Zhang et al. (Zhang et al., 2015) has traditionally been applied in offline CL settings. Recent advancements have adapted this benchmark for online CL, notably in studies such as MBPA++ (de Masson D’Autume et al., 2019) and OML-ER (Holla et al., 2020)
@@ 7.3. Offline CL Datasets for Multi-modal Tasks
LAION-400M (Schuhmann et al., 2021) and LAION-5B (Schuhmann et al., 2022) are both CLIP-filtered, large-scale, open-access datasets designed to support multimodal models like CLIP, but they differ significantly in scale and scope. LAION-400M contains 400 million image-text pairs primarily in English, while LAION-5B is a much larger and more diverse dataset with 5.85 billion pairs, including multilingual and language-agnostic examples. CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021) are large-scale datasets for vision-and-language pre-training, differing in scale, filtering, and diversity. CC3M has 3.3 million image-text pairs with strict filtering for high-precision captions, while CC12M, with 12.4 million pairs, relaxes these filters to capture a broader range of concepts, especially long-tail categories. COYO-700M (Byeon et al., 2022) contains 747 million image-text pairs, designed to train various models with additional meta-attributes. Datacomp (Gadre et al., 2024) is a benchmark designed to advance the creation of multimodal datasets, particularly for image-text models such as CLIP.
@@ 8. Metrics
Memory stability is typically evaluated by backward transfer (Lopez-Paz and Ranzato, 2017) and forgetting measure (Chaudhry et al., 2018). Backward Transfer (BWT) emerges as a pivotal concept extensively documented in the literature, notably by Lopez et al. (Lopez-Paz and Ranzato, 2017) and Wu et al. (Wu et al., 2023). BWT measures the performance degradation on previously mastered tasks after the model is trained on new tasks. This performance degradation phenomenon is often referred to as “forgetting”.
@@
Additionally, Chaudhry et al. (Chaudhry et al., 2018) introduce the Forgetting Measure (FM), a metric designed to quantify the extent of forgetting a model experiences for a specific task. A lower FM indicates better retention of previous tasks. Davari et al. (Davari et al., 2022) propose linear probes (LP) to assess representation forgetting via an optimal linear classifier trained on the frozen activations of a base network. Representation forgetting is quantified by evaluating the change in Language Processing (LP) performance before and after the introduction of a new task. Kemker et al. (Kemker et al., 2018) introduce three metrics, where $\Omega_{\text{base}}$ assesses retention of initial learning, $\Omega_{\text{new}}$ measures recall of new tasks, and $\Omega_{\text{all}}$ evaluates overall proficiency in maintaining old knowledge and acquiring new information. Additionally, researchers (Koh et al., 2022) devise a novel metric, termed the Knowledge Loss Ratio (KLR), quantifies knowledge degradation using principles from information theory.
@@
Evaluating learning plasticity can be effectively accomplished through two key metrics: forward transfer (FWT) (Lopez-Paz and Ranzato, 2017) and intransigence measure (IM) (Chaudhry et al., 2018). Forward Transfer (FWT) (Lopez-Paz and Ranzato, 2017) assesses the beneficial effects on the performance of subsequent tasks following a model’s training on prior tasks.
@@
where $R_{0,i}$ denotes the performance metric associated with training on task $i$ independently. Higher values of FWT indicate superior model performance.
Intransigence measure (IM), as defined by Chaudhry et al. (Chaudhry et al., 2018), quantifies a model’s inability to learn new tasks by comparing the performance difference of a task when trained jointly with other tasks versus. Moreover, Koh et al. (Koh et al., 2022) introduce novel metrics, known as Knowledge Gain Ratio (KGR), which quantifies the capacity to acquire new knowledge by calculating knowledge gain.
Table 2. Comparison of typical continual learning methods, where Reg. and Para. mean Regularization and Parameter-isolation.
CL settings Method FM Arch. Continual Learning Tech.
Traditional Continual Pre-training Parameter-Efficient Instruction Tuning
DIL LFPT5 (Qin and Joty, 2022b) T5 Replay — Reg. — Para.
B-CL (Ke and Xu, 2021) BERT Para.
ELLE (Qin et al., 2022) BERT — GPT Reply — Para.
AdapterCL (Madotto et al., 2020) GPT2 Para.
DEMIX (Gururangan et al., 2022) GPT2 — GPT3 Para.
CLASSIC (Ke et al., 2021b) BERT Para.
LLM-CL (Ding et al., 2024) LLaMA Replay — Reg. — Para.
CPT (Ke et al., 2022) RoBERTa Para.
C-PT (Zhu et al., 2022) T5 Para.
CL-KD (Castellucci et al., 2021) BERT Reg.
PlugLM (Cheng et al., 2023b) BERT Para.
Pretr (Cossu et al., 2022) BERT — RoBERTa Para.
AEWC (Lee, 2017) LSTM-RNNs Reg.
Continual DAP-training (Ke et al., 2023) RoBERTa Reg.
COPF (Zhang et al., 2023b) LLaMA Replay — Reg. — Para.
LAMOL (Sun et al., 2020a) GPT2 Replay
RVAE_LAMOL (Wang et al., 2022a) GPT2 Replay
Adapt-Retrieve-Revise (Zhang et al., 2023c) Baichuan RAG
Lifelong-MoE (Chen et al., 2023) GLaM Para.
CPPO (Zhang et al., 2024b) GPT2 Reg.
EcomGPT-CT (Ma et al., 2023) BLOOM Replay
S-Prompt (Wang et al., 2022b) CLIP Para.
VQACL (Zhang et al., 2023d) VL-T5 Replay — Reg.
TIL PP (Razdaibiedina et al., 2023) BERT — T5 Para.
CTR (Ke et al., 2021a) BERT Para.
MeLL (Wang et al., 2021) BERT — RoBERTa Replay
ERDA (Qin and Joty, 2022a) Bi-LSTM — BERT Replay — Reg.
PCLL (Zhao et al., 2022) GPT2 Replay
BiHNet-Reg (Jin et al., 2021) BART — BART-Adapter Para.
ConTinTin (Yin et al., 2022) BART Replay — InstructionSpeak
HMI (Maekawa et al., 2023) BERT, GPT2 Replay
ACM (Zhang et al., 2022) GPT2 Replay — Para.
DYNAINST (Mok et al., 2023) BART Replay — Reg.
Conure (Yuan et al., 2021) TCN Para.
TERACON (Kim et al., 2023) NextitNet Para.
ERNIE 2.0 (Sun et al., 2020b) ERNIE Replay
RecyclableTuning (Qin et al., 2023) RoBERTa Replay — Para.
Conpet (Song et al., 2023) LLaMA Replay — Para.
InstructAlign (Cahyawijaya et al., 2023) BLOOMZ Replay
Continual-T0 (Scialom et al., 2022b) T0 Replay
DynaMind (Du et al., 2023) LLaMA — GPT — Falcon Replay — Para.
ELM (Jang et al., 2023) T5 Para.
O-LoRA (Wang et al., 2023a) LLaMA — Alpaca Para.
COPF (Zhang et al., 2023b) LLaMA Replay — Reg. — Para.
SLM (PENG et al., 2024) LLaMA, T5, BERT Para.
CTP (Zhu et al., 2023) BERT — ViT Para. -
ZSCL (Zheng et al., 2023) CLIP Replay — Reg.
MoE-Adapters4CL (Yu et al., 2024b) CLIP Para.
TRIPLET (Qian et al., 2023) ALBEF — FLAVA Para.
RecAdam (Chen et al., 2020) BERT — ALBERT Reg.
CIL EPI (Wang et al., 2023c) BERT Para.
IDBR (Huang et al., 2021) BERT Replay
PAGeR (Varshney et al., 2022) GPT2 Replay
ENTAILMENT (Xia et al., 2021) RoBERTa prompt-based
ExtendNER (Monaikul et al., 2021) BERT Reg.
PLE (Li et al., 2022b) RoBERTa Replay — Para.
DE&E (Wójcik et al., 2023) Distilbert Para.
SRC (Liu et al., 2019b) SIF Reg.
MoE-Adapters4CL (Yu et al., 2024b) TIL RandSel — KMeansSel
VLM-PL (Kim et al., 2024) CLIP — Ferret Replay
Adaptation-CLIP (Liu et al., 2023a) CLIP Para.
PROOF (Zhou et al., 2023c) CLIP Para.
LGCL (Khan et al., 2023) CLIP — ViT Para.
ZSCL (Zheng et al., 2023) CLIP Replay — Reg.
CLAP (Jha et al., 2024) CLIP Replay — Para.
GMM (Cao et al., 2024) EVA-CLIP — BLIP2 Reg.
Online CL MBPA++ (de Masson D’Autume et al., 2019) BERT Replay
Meta-MBPA++ (Wang et al., 2020) BERT Replay
OML-ER (Holla et al., 2020) BERT Replay
TPEM (Geng et al., 2021) GRU Para.
MSR (Liu et al., 2021b) BERT Replay — Reg.
ProgModel (Shen et al., 2019) RNN Para.
DKR (Cui et al., 2024) CLIP Reg.
GMM (Cao et al., 2024) EVA-CLIP — BLIP2 — MiniGPT-4 Para.
@@
CKL (Jang et al., 2022b) introduces FUAR, which quantitatively measures the efficiency of each CKL method. It calculates the instances of time-invariant knowledge that a model forgets to learn or update one instance of new knowledge.
Table 3. A summary of the performance of various Foundation LMs. A-RG: Average of ROUGE-1, -2 and -L scores; F1: F1 score; Acc.: Accuracy; MF1: Macro-F1; AMF1: averaged Macro-F1; JGA: Joint Goal Accuracy; SER: Slot Error Rate; AA: average accuracy; AIA: average incremental accuracy; FM: forgetting measure.
Methods Dataset F1 Acc A-RG MF1 JGA SER BLEU AA AIA FM FWT BWT AMF1
LFPT5 (Qin and Joty, 2022b) AGNews, Amazon, DBPedia, Yahoo - 52.71 - - - - - - - - - - -
SLM(T5) (PENG et al., 2024) AGNews, Amazon, DBpedia, Yahoo - - - - - - - 73.10 - - - - -
PP(T5) (Razdaibiedina et al., 2023) AGNews, Amazon, DBpedia, Yahoo - - - - - - - 75.10 - - - - -
MBPA++ (de Masson D’Autume et al., 2019) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 70.60 - - - - -
IDBR (Huang et al., 2021) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 73.19 - - - - -
EPI (Wang et al., 2023c) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 74.43 - - - - -
OML-ER (Holla et al., 2020) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 75.70 - - - - -
O-LoRA(Wang et al., 2023a) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 75.80 - - - - -
LAMOL (Sun et al., 2020a) AGNews, Amazon, DBPedia, Yahoo, Yelp - - - - - - - 76.50 - - - - -
PP(Bert) (Razdaibiedina et al., 2023) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 77.90 - - - - -
Meta-MBPA++ (Wang et al., 2020) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 77.30 - - - - -
SLM(Bert) (PENG et al., 2024) AGNews, Amazon, DBpedia, Yahoo, Yelp - - - - - - - 79.10 - - - - -
B-CL (Ke and Xu, 2021) HL5Domains, Liu3Domains, Ding9Domains, SemEval14 - 88.29 - 81.40 - - - - - - - - -
CLASSIC (Ke et al., 2021b) HL5Domains, Liu3Domains, Ding9Domains, SemEval14 - 90.22 - 85.12 - - - - - - - - -
CTR (Ke et al., 2021a) HL5Domains, Liu3Domains, Ding9Domains, SemEval14 - 89.47 - 83.62 - - - - - - - - -
LLM-CT (Ding et al., 2024) HL5Domains, Liu3Domains, Ding9Domains, SemEval14 - 94.91 - 91.43 - - - - - - - - -
PP(Bert) (Razdaibiedina et al., 2023) MNLI, CB, WiC, COPA, QQP, BoolQ, RTE, IMDB, Yelp, Amazon, SST2, DBpedia, AGNews, Yelp, Amazon, Yahoo - - - - - - - 69.3 - - - - -
O-LoRA(Wang et al., 2023a) MNLI, CB, WiC, COPA, QQP, BoolQ, RTE, IMDB, Yelp, Amazon, SST2, DBpedia, AGNews, Yelp, Amazon, Yahoo - - - - - - - 69.6 - - - - -
PP(T5) (Razdaibiedina et al., 2023) MNLI, CB, WiC, COPA, QQP, BoolQ, RTE, IMDB, Yelp, Amazon, SST2, DBpedia, AGNews, Yelp, Amazon, Yahoo - - - - - - - 79.5 - - - - -
OML-ER (Holla et al., 2020) FewRel - 69.5 - - - - - - - - - - -
DYNAMIC CONPET(Song et al., 2023) FewRel - 88.62 - - - - - 88.62 - - - - -
PAGeR (Varshney et al., 2022) FewRel - - - - - - - 91.30 - - - - -
CPPO (Zhang et al., 2024b) IMDB, HH-RLHF, Reddit TL, DR - - - - - - - 77.40 84.20 2.9 - -2.9 -
COPF (Zhang et al., 2023b) IMDB, HH-RLHF, Reddit TL, DR - - - - - - - 86.30 85.10 -0.60 - 0.60 -
LAMOL (Sun et al., 2020a) SST, QA-SRL, WOZ - - - - - - - 81.00 - - - - -
RVAE_LAMOL (Wang et al., 2022a) SST, QA-SRL, WOZ - - - - - - - 81.60 - - - - -
PLE(5-shot regime) (Li et al., 2022b) CLINC150, ATIS, HWU64, BANKING77, MTOP, SNIPS, LEYZER, MSLU, TOP - - - - - - - 85.61 - - - - -
PLE(10-shot regime) (Li et al., 2022b) CLINC150, ATIS, HWU64, BANKING77, MTOP, SNIPS, LEYZER, MSLU, TOP - - - - - - - 88.16 - - - - -
MBPA++ (de Masson D’Autume et al., 2019) QuAC, TrWeb, TrWik, SQuAD - - - - - - - - - - - - 62.40
Meta-MBPA++ (Wang et al., 2020) QuAC, TrWeb, TrWik, SQuAD - - - - - - - - - - - - 64.90
CPT (Ke et al., 2022) Yelp Restaurant, AI Papers, ACL Papers, AGNews - 52.59 - 46.41 - - - - - - - - -
Continual DAP-training (Ke et al., 2023) Yelp Restaurant, Amazon Phone, Amazon Camera, ACL Papers, AI Papers, PubMed Papers - 81.91 - 77.93 - - - - - - - - -
DYNAMIC CONPET(Song et al., 2023) OntoNotes - 84.47 - - - - - 85.83 - - - - -
LFPT5 (Qin and Joty, 2022b) CoNLL03, OntoNotes 47.59 - - - - - - - - - - - -
AdapterCL (Madotto et al., 2020) TM19, TM20, SGD, MultiWoZ - 90.50 - - 35.10 31.78 16.76 - - - - - -
RMR_ DSE (Li et al., 2022a) MultiWoZ-2.0 - - - - - 48.79 39.86 - - - - - -
DEMIX (Gururangan et al., 2022) In-Car Assistant, Multi-WOZ 2.1, CamRest 45.84 - - - - - 11.96 - - - - - -
LFPT5 (Qin and Joty, 2022b) CNNDM, WikiHow, XSum - - 17.05 - - - - - - - - - -
C-PT (Zhu et al., 2022) Schema-Guided Dialog dataset (SGD) - - - - 61.20 - - - - - 13.70 0.50 -
Lifelong-MoE (Chen et al., 2023) TriviaQA 20.22 - - - - - - - - - - - -
Lifelong-MoE (Chen et al., 2023) WMT16 - - - - - - 19.16 - - - - - -
CTR (Ke et al., 2021a) 10 DSC datasets - 89.31 - 88.75 - - - - - - - - -
CTR (Ke et al., 2021a) 20News - 95.25 - 95.23 - - - - - - - - -
MeLL (Wang et al., 2021) TaskDialog-EUIC - - - - - - - 93.79 - - - - 93.42
MeLL (Wang et al., 2021) Hotline-EUIC - - - - - - - 96.73 - - - - 93.41
PCLL (Zhao et al., 2022) HWU, BANKING, CLINC, SNIPS, TOP - - - - - - - 90.25 - - - - -
PCLL (Zhao et al., 2022) SNIPS, AITS, DSTC, MIT-MOVIE, MIT-RESTAURANT - - - - - - - - - - - - 74.48
SLM(LLaMA) (PENG et al., 2024) Medical, MMLU, Finance - - - - - - - 82.3 - - - - -
EPI (Wang et al., 2023c) Web of Science (WOS) - - - - - - - 77.83 - - - - -
IDBR (Huang et al., 2021) Yahoo, AGNews, Yelp - - - - - - - 72.53 - - - - -
@@
Near-future accuracy (NFA) (Al Kader Hammoud et al., 2023) is introduced as a novel evaluation metric for OCL problems. Unlike traditional evaluation methods that assess models on immediately subsequent samples, NFA evaluates models on samples slightly further into the future, using a minimal shift $S$. Such operation can mitigate label correlation effects, which can adversely impact the accuracy of model adaptability assessments. Yogatama et al. (Yogatama et al., 2019) proposed a novel online codelength, inspired by prequential encoding (Blier and Ollivier, 2018), to quantify how quickly an existing model can adapt to a new task.
9. Analysis and Discussion
In this section, we compare the performance and characteristics of typical foundation model-based continual learning methods (Table 2 and Table 3).
The effectiveness of each continual learning method is shaped by the specific task environment, resource constraints, and application goals. Replay-based methods like LAMOL, Meta-MBPA++, PAGeR, and COPF excel in scenarios where memory efficiency and maintaining a balance between old and new knowledge are essential. These methods strike a delicate balance between retaining previously acquired knowledge and adapting to new tasks. Parameter isolation is key to minimizing interference between highly distinct tasks. By allocating dedicated parameters to each task, these approaches prevent overlap and ensure that task-specific learning is preserved. Techniques such as AdapterCL, B-CL, and EPI are particularly effective in environments with highly variable tasks, where reducing interference is critical. Regularization techniques, on the other hand, enhance stability in structured task settings. These methods constrain model updates to protect knowledge from previous tasks, making them ideal in situations with well-defined task boundaries. Regularization-based approaches like RMR_DSE, CLASSIC, and EWC mitigate catastrophic forgetting by limiting model changes, thereby preserving prior learning. These methods are particularly effective in domain-incremental learning, where tasks evolve gradually, and long-term stability is vital.
First, most continual learning (CL) methods are built on PLMs like BERT, RoBERTa, and BART, or smaller-scale LLMs such as LLaMA. Models leveraging LLMs often outperform PLM-based ones, thanks to their greater capacity for handling complex tasks and enhanced adaptability. Second, the majority of CL methods are tailored for offline settings, where the full target dataset is assumed to be available for training. However, in real-world scenarios, data typically arrives in a streaming fashion, making it challenging to anticipate distribution shifts and effectively manage dynamic data. Third, a significant hurdle in the field is the lack of unified benchmarks and standardized metrics for evaluating CL algorithms. This fragmentation complicates the comparison of methods across tasks and domains. To propel the field forward, there is a critical need for benchmark datasets and robust metrics that can consistently gauge the performance of CL models, particularly in dynamic, data-streaming environments. These benchmarks should span diverse tasks and modalities to ensure comprehensive evaluation across various domains.
@@ 10. Challenges and Further Work
Most existing studies in the domain of continual learning assume static datasets with known distributions in a relatively closed environment. Moreover, these studies mainly focus on simple tasks (e.g., text classification, sentiment analysis and intent classification) with clear labels. These assumptions do not hold in real-world applications, where environments continually evolve and introduce novel stimuli. A key challenge is developing continual learning models that can autonomously detect and adapt to data distribution shifts, thereby improving the applicability of AI in dynamic real-world scenarios. Liu et al. (Liu et al., 2023b) recently proposed the SOLA framework to address these limitations by facilitating autonomous adaptation in AI systems. Despite this progress, significant challenges remain in enabling these systems to adjust to new, dynamic environments without ongoing human oversight.
Bridging Machine Learning and Cognitive Science in Continual Learning.
While continual learning aims to mimic human-like learning, a gap persists in how well these methods align with cognitive science findings on human continuous learning. In human cognition, processes such as rehearsal, memory consolidation, and adaptive forgetting are key to learning in dynamic environments. In contrast, machine learning models often face challenges like catastrophic forgetting, a phenomenon without a clear biological equivalent. Bridging these fields could enhance the design of continual learning algorithms by drawing on cognitive science insights, potentially leading to more robust, adaptive, and human-like learning systems.
@@
Privacy protection in continual learning systems poses a significant challenge, particularly as these systems are designed to continuously update and refine their models based on incoming data streams. Unlike traditional static machine learning models, continual learning systems frequently access and process sensitive data across different contexts and time periods, raising substantial concerns about data confidentiality and user privacy. Effective privacy-preserving mechanisms must be integrated into the architecture of these systems to ensure that they do not inadvertently expose or misuse personal data. Techniques such as differential privacy (Dwork, 2006), federated learning (Zhang et al., 2021), and secure multi-party computation (Goldreich, 1998) offer promising solutions by allowing models to learn from decentralized data sources without needing to access the actual data directly.
Continual Alignment.
In dynamic real-world environments, human preferences and requirements are ever-evolving, making continual alignment crucial for maintaining AI system effectiveness. Traditional alignment methods often assume static preferences, which is insufficient in changing contexts. Integrating Reinforcement Learning from Human Feedback (RLHF) into continual learning presents a unique challenge, as models must adapt to new preferences while retaining previously learned ones. Future research should prioritize developing methods like CPPO and COPF that balance integrating new feedback with preserving past alignment, ensuring long-term adaptability to human needs.